# Are we there yet? A brief survey of Music Emotion Prediction Datasets, Models and Outstanding Challenges

### v1

@@ Abstract
Deep learning models for music have advanced drastically in the last few years. But how good are machine learning models at capturing emotion these days and what challenges are researchers facing? In this paper, we provide a comprehensive overview of the available music-emotion datasets and discuss evaluation standards as well as competitions in the field. We also provide a brief overview of various types of music emotion prediction models that have been built over the years, offering insights into the diverse approaches within the field. Through this examination, we highlight the challenges that persist in accurately capturing emotion in music. Recognizing the dynamic nature of this field, we have complemented our findings with an accompanying GitHub repository. This repository contains a comprehensive list of music emotion datasets and recent predictive models [^1] https://github.com/AMAAI-Lab/awesome-MER/.
1 Introduction
@@
With the advent of technology and data-driven methodologies, new avenues have opened up for exploring the complex relationship between music and emotions, such as deep learning models. Despite these advancements, accurate predictions of emotions from music remains elusive. This is due to a number of reasons, including the subjective, personal nature of emotional perception, as well as bias and limitations in the current datasets, and challenges in benchmarking. We discuss these and other challenges at length in Section 5.
Despite these challenges, the potential applications of Music Emotion Recognition (MER) systems are vast and varied. In the healthcare domain, we may find personalized music recommendation systems that guide a listener to different emotional states [5], or we may even see MER systems used to build large-scale datasets on which we can train generative music AI systems that can be controlled to generate music with specific emotions [6, 7]. In the same line of thought, [8] developed a Brain-Computer Interface system that can provide musical feedback about the listener’s current emotional state and subsequently influence this state. Additionally, MER systems may help facilitate emotional analysis during music composition, interactive experiences in media and entertainment, as well as inform market research [9]. The implications of understanding music and emotions extend across diverse domains. In general, following the idea of positive psychology [10], understanding music emotions can be used to enhance the user experience when designing various systems.
@@
In the next section, we provide an extensive overview of the available emotion-annotated music datasets. This is followed by a discussion of the evaluation practices in the field (Section 3). After that, we describe a selected list of recent models and approaches in Section 4. Finally, Section 5 dives into the remaining challenges and future direction for the field of MER, followed by a general conclusion.
@@
Table 1: Overview of Music Emotion Datasets (Sorted by Year).
Static/ Perceived/
Dataset Year # of instances Length Type Categorical Dimensional Dynamic Induced
MoodsMIREX [15] 2007 269 30s MP3 5 labels - Static Perceived
CAL500 [16] 2007 500 full MP3 174 labels - Static Perceived
Yang-Dim [17] 2008 195 25s WAV - Russell Static Perceived
MoodSwings [18] 2008 240 15s MP3 - Russell Dynamic Perceived
NTWICM [19] 2010 2,648 full MP3 - Russell Static Perceived
Soundtracks [20] 2011 470 15s-1m MP3 6 labels 3 dimemsions Static Perceived
DEAP [21] 2012 120 60s YouTube id - Russell Static Induced
Panda et al.’s dataset [22] 2013 903 30s MP3, MIDI 21 labels - Static Perceived
Solymani et al.’s dataset [23] 2013 1000 45s MP3 - Russell Both Perceived
Emotify [24] 2016 400 60s MP3 GEMS - Static Induced
Moodo [25] 2016 200 15s WAV - Russell Static Perceived
CH818 [26] 2017 818 30s MP3 - Russell Static Perceived
4Q-emotion [27] 2018 900 30s MP3 Quadrants - Static Perceived
MediaEval DEAM [28] 2018 2,058 45s MP3 - Russell Both Perceived
PMEmo [29] 2018 794 full MP3 - Russell Both Induced
RAVDESS [30] 2018 1,012 full MP3, MP4 5 labels - Static Perceived
DMDD [31] 2018 18,644 full Audio, Lyrics - Russell Static Perceived
MTG-Jamendo [32] 2019 18,486 full MP3 56 labels - Static Perceived
VGMIDI [33] 2019 200 full MIDI - Russell Dynamic Perceived
Turkish Music Emotion [34] 2019 400 30s MP3 4 labels - Static Perceived
EMOPIA [35] 2021 1,087 30s-40s Audio, MIDI Quadrants - Static Perceived
MER500 [36] 2020 494 10s WAV 5 labels - Static Perceived
Music4all [37] 2020 109,269 30s WAV - 3 dimensions Static Perceived
CCMED-WCMED [38] 2020 800 8-20s WAV - Russell Static Perceived
MuSe [39] 2021 90,001 full Audio - Russell (V-A-D) Static Perceived
HKU956 [40] 2022 956 full MP3 - Russell Static Induced
MERP [41] 2022 54 full WAV - Russell Both Perceived
MuVi [42] 2022 81 full YouTube id GEMS Russell Both Perceived
YM2413-MDB [43] 2022 699 full WAV, MIDI 35 labels - Static Perceived
MusAV [44] 2022 2,092 full WAV Russell Static Perceived
EmoMV [45] 2023 5,986 30s WAV 6 labels - Static Perceived
Indonesian Song [46] 2023 476 full WAV 3 labels - Static Perceived
TROMPA-MER [47] 2023 1,161 30s WAV 11 labels - Static Perceived
EMMA [48] 2024 364 30s-60s WAV GEMS - Static Perceived
SiTunes [49] 2024 300 full WAV - Russell Static Induced
2 Datasets
@@
Table 1 presents an extensive overview of emotion-annotated music datasets. These datasets vary in size, annotation granularity, and focus on either perceived or induced emotions. Before delving deeper into the datasets, it is essential to understand the emotion representations commonly referenced in the field.
@@
Categorical models include the Geneva Emotional Music Scales (GEMS) [24]. This model was specifically designed for music-induced emotions and consists of 45 emotion tags grouped into nine categories, including amazement, solemnity, tenderness, nostalgia, calmness, power, joyful activation, tension, and sadness. The datasets listed in Table 1 use a variety of emotion representation models, with Russell’s model being the most popular dimensional representation. Some datasets do not use a specific emotion model, for instance, the MTG-Jamendo dataset consists of freely assigned tags by the listeners, resulting in a diverse and comprehensive set of 56 tags, spanning from ‘melancholic’ to ‘upbeat’. A subset of this dataset is used for the ‘Emotion and Theme Recognition in Music’ Task of MediaEval. In Table 1, the number of tags used to represent the emotions are listed in the column ‘Categorical’.
@@
Induced versus perceived Emotion labels in the datasets can represent perceived or induced emotion. Perceived emotions are emotions that listeners consciously recognize while listening to the music. Induced emotions, on the other hand, are emotions that listeners experience as a result of listening to the music, but are not necessarily consciously recognized. Most of the datasets in Table 1 use perceived emotion labels, which are easier to annotate. For induced emotion labels, researchers have used psycho-physiological measurements ranging from electromyogram (EMG), volume pulse (BVP), electrocardiograms (ECG), skin conductance, respiration rate, heart rate, to electroencephalograms (EEG). There have been many studies in psychology that use such biosensors to explore how music can influence our emotions [54, 55, 56, 57, 58]. Since these studies include medical data, the datasets are not often public. However, there are a few datasets with music and its induced emotions. Firstly the DEAP dataset [21] includes EEG, facial video recordings, as well as peripheral physiological signals that were recorded while they watched music videos. In addition, they collected perceived emotion ratings in terms of arousal, valence, like/dislike, dominance and familiarity. Second, the HKU956 dataset [40] records five kinds of physiological signals (i.e., heart rate, electrodermal activity, blood volume pulse, inter-beat interval, and skin temperature) of participants as they listen to music, along with reported emotions in the arousal and valence dimensions. Finally, SiTunes [49] includes physiological signals (i.e., heart rate, activity intensity, activity step, and activity type) measured both before and after music listening, alongside environmental data (i.e., time of day, weather information, and location) recorded during users’ daily lives. The MER models that can predict induced emotions have various medical applications, such as curating playlists to guide patients to different emotional states [5]. An interesting study by [59] concludes that music tends to induce the emotion that is perceived, enabling researchers to use perceived emotion datasets for developing emotion-inducing models.
@@
Modalities Music comes in many formats, the most common one being ‘audio’. Audio files can either be raw waveforms or compressed .mp3 files. However, we should not neglect the MIDI format, a popular format still often used by music producers, composers and performers. Whereas most datasets contain audio files, only a handful focus on MIDI: 1) the VGMIDI dataset [33], which contains continuous valence/arousal ratings for MIDI files of piano arrangements for video game soundtracks; 2) the Panda et al.’s dataset [22] (used for the MIREX competition), which provides a diverse collection of audio clips, lyrics, and aligned MIDI files, contains labels for five emotion clusters derived from a cluster analysis of online tags; and 3) the EMOPIA dataset [35], which contains paired piano music audio with MIDI that has emotion annotations of the four high/low valence/arousal quadrants.
@@
Many sources can elicit emotion. For instance, video, or lyrics may also affect our perceived or induced emotions. Some datasets offer alternative multimedia streams such as emotion-rated music videos from a variety of genres, including pop, rock, classical, and jazz as in the DEAP dataset [21]; or text of lyrics with the annotated music in the DMDD dataset [31]. Finally, the MuVi dataset [42] even offers isolated modality ratings for music videos. In this study, the raters were presented with either the music video, the music alone, or the muted video. The final dataset contains ratings for each of these modalities separately as well as together. This allows [42] to build a model on pure isolated modalities, which proved to be more accurate than a traditional model.
@@
Dataset size Compared to affective datasets available in other domains (e.g., the Sentiment140 dataset [60] in NLP, which contains 1.6 million tweets annotated as positive or negative), the size of the available datasets is still very limited, with the largest dataset containing 109,269 instances. In addition to the number of instances in datasets, we also notice a difference in the length of the instances. A number of datasets (e.g. MERP [41], VGMIDI [33], and HKU956 [40]) offer ratings for full-length songs. In datasets with dynamic ratings throughout the song, this may provide a means for researchers to analyse how our emotions evolve throughout a song. Other datasets focus on short fragments, often ranging from 30 seconds to 1-minute fragments (e.g. DEAM [28], EMOPIA [35], and EMMA [48]), but even as short as 10s as is the case for the MER500 dataset [36].
@@
In sum, there is a variety of emotion-annotated datasets available as shown in Table 1. They differ in terms of emotion representation models, as well as music format and perceived versus induced emotion labels. The challenges and future research directions related to datasets are discussed in detail in Section 5.
3 Evaluation Protocols
@@
Evaluation metrics play a crucial role in assessing the performance of MER systems. Depending on which type of emotion ratings are being predicted (dimensional versus categorical), the evaluation metrics change as we are dealing with a regression or a classification task respectively. Commonly used evaluation metrics for categorical MER systems include accuracy, precision, Area under the Curve (AUC), and confusion matrices. In the case of regression MER models, we see metrics such as precision, recall, F1-score, mean squared error (MSE), $R^{2}$, and Pearson correlation coefficient [61]. These metrics provide insights into the effectiveness of MER models when it comes to emotions represented by dimensional models.
@@
Evaluating MER systems goes beyond just defining a common metric. Due to the inherent differences between datasets, a comparison across datasets is often not possible. For an in-depth discussion on this, the reader is referred to Section 5. Within one dataset, however, it is possible to establish benchmarks and compare the performance of different models, provided that the train/test split is shared. There are some initiatives to facilitate the comparison between models, such as competitions, as well as individual papers that offer clear data splits and metrics [62, 63, 64, 65, 66].
@@
Competitions and challenges provide valuable platforms for evaluating and comparing the performance of MER systems. These initiatives often involve standardized datasets, evaluation protocols, and metrics, enabling researchers to benchmark their algorithms against state-of-the-art methods. Existing benchmarking initiatives and competitions in MER include the ‘Audio K-POP Mood Classification’ task in MIREX (Music Information Retrieval Evaluation eXchange) (last organized in 2019) [^2] https://www.music-ir.org/mirex/wiki/2019:Audio_K-POP_Mood_Classification/ , the ‘Emotion in Music’ task in MediaEval (last organized in 2015) [^3] http://www.multimediaeval.org/mediaeval2015/emotioninmusic2015/, and the ‘Emotion and Theme Recognition in Music Using Jamendo’ in MediaEval (last organized in 2021) [^4] https://multimediaeval.github.io/2021-Emotion-and-Theme-Recognition-in-Music-Task/.
@@
4 Models and approaches
In this section, we briefly touch upon some existing MER models. This is by no means a complete overview, for this, the reader is referred to other survey papers [11, 12, 13, 9, 14]. The aim of this section is merely to point out the current state-of-the-art and various approaches over the last few years.

### v2

@@ Abstract
Deep learning models for music have advanced drastically in recent years, but how good are machine learning models at capturing emotion, and what challenges are researchers facing? In this paper, we provide a comprehensive overview of the available music-emotion datasets and discuss evaluation standards as well as competitions in the field. We also offer a brief overview of various types of music emotion prediction models that have been built over the years, providing insights into the diverse approaches within the field. Through this examination, we highlight the challenges that persist in accurately capturing emotion in music, including issues related to dataset quality, annotation consistency, and model generalization. Additionally, we explore the impact of different modalities, such as audio, MIDI, and physiological signals, on the effectiveness of emotion prediction models. Recognizing the dynamic nature of this field, we have complemented our findings with an accompanying GitHub repository111 https://github.com/AMAAI-Lab/awesome-MER/. This repository contains a comprehensive list of music emotion datasets and recent predictive models.
Index Terms:
Deep learning, artificial intelligence, music emotion recognition.
I Introduction
@@
With the advent of technology and data-driven methodologies, new avenues have opened up for exploring the complex relationship between music and emotions, such as deep learning models. Despite these advancements, accurate predictions of emotions from music remain elusive. This is due to a number of reasons, including the subjective, personal nature of emotional perception, as well as bias and limitations in the current datasets, and challenges in benchmarking. We discuss these and other challenges at length in Section V.
Despite these challenges, the potential applications of Music Emotion Recognition (MER) systems are vast and varied. In the healthcare domain, we may find personalized music recommendation systems that guide a listener to different emotional states [5], or we may even see MER systems used to build large-scale datasets on which we can train generative music AI systems that can be controlled to generate music with specific emotions [6, 7, dashherremans2017morpheus]. In the same line of thought, [8] developed a Brain-Computer Interface system that can provide musical feedback about the listener’s current emotional state and subsequently influence this state. Additionally, MER systems may help facilitate emotional analysis during music composition, interactive experiences in media and entertainment, as well as inform market research [9]. The implications of understanding music and emotions extend across diverse domains. In general, following the idea of positive psychology [10], understanding music emotions can be used to enhance the user experience when designing various systems.
@@
In the next section, we provide an extensive overview of the available emotion-annotated music datasets. This is followed by a discussion of the evaluation practices in the field (Section III). After that, we describe a selected list of recent models and approaches in Section IV. Finally, Section V dives into the remaining challenges and future direction for the field of music emotion recognition (MER), followed by a general conclusion.
@@
TABLE I: Overview of Music Emotion Datasets (Sorted by Year). Note: S = Static, D = Dynamic, B = Both; P = Perceived, I = Induced.
Dataset Year # of instances Length Type Categorical Dimensional S/D/B P/I
MoodsMIREX [15] 2007 269 30s MP3 5 labels - S P
CAL500 [16] 2007 500 full MP3 174 labels - S P
Yang-Dim [17] 2008 195 25s WAV - Russell S P
MoodSwings [18] 2008 240 15s MP3 - Russell D P
NTWICM [19] 2010 2,648 full MP3 - Russell S P
Soundtrack [20] 2011 470 15s-1m MP3 6 labels 3 dimensions S P
MoodSwings Turk [21] 2011 240 15s MP3 - Russell D P
DEAP [22] 2012 120 60s YouTube id - Russell S I
Panda et al.’s dataset [23] 2013 903 30s MP3, MIDI 21 labels - S P
Solymani et al.’s dataset [24] 2013 1,000 45s MP3 - Russell B P
CAL500exp [25] 2014 3,223 3s-16s MP3 67 labels - S P
AMG1608 [26] 2015 1,608 30s WAV - Russell S P
Emotify [27] 2016 400 60s MP3 GEMS - S I
Moodo [28] 2016 200 15s WAV - Russell S P
Malheiro et al.’s dataset [29] 2016 200 30s Audio, Lyrics Quadrants - S P
CH818 [30] 2017 818 30s MP3 - Russell S P
4Q-emotion [31] 2018 900 30s MP3 Quadrants - S P
DEAM [32] 2018 2,058 45s MP3 - Russell B P
PMEmo [33] 2018 794 full MP3 - Russell B I
RAVDESS [34] 2018 1,012 full MP3, MP4 5 labels - S P
DMDD [35] 2018 18,644 full Audio, Lyrics - Russell S P
MTG-Jamendo [36] 2019 18,486 full MP3 56 labels - S P
VGMIDI [37] 2019 200 full MIDI - Russell D P
Turkish Music Emotion [38] 2019 400 30s MP3 4 labels - S P
EMOPIA [39] 2021 1,087 30s-40s Audio, MIDI Quadrants - S P
MER500 [40] 2020 494 10s WAV 5 labels - S P
Music4all [41] 2020 109,269 30s WAV - 3 dimensions S P
CCMED-WCMED [42] 2020 800 8-20s WAV - Russell S P
MuSe [43] 2021 90,001 full Audio - Russell (V-A-D) S P
HKU956 [44] 2022 956 full MP3 - Russell S I
MERP [45] 2022 54 full WAV - Russell B P
MuVi [46] 2022 81 full YouTube id GEMS Russell B P
YM2413-MDB [47] 2022 699 full WAV, MIDI 35 labels - S P
MusAV [48] 2022 2,092 full WAV - Russell S P
EmoMV [49] 2023 5,986 30s WAV 6 labels - S P
Indonesian Song [50] 2023 476 full WAV 3 labels - S P
TROMPA-MER [51] 2023 1,161 30s WAV 11 labels - S P
Music-Mouv [52] 2023 188 full Spotify id GEMS - S I
ENSA [53] 2023 60 full MP3 - Russell D P
EMMA [54] 2024 364 30s-60s WAV GEMS - S P
SiTunes [55] 2024 300 full WAV - Russell S I
MERGE [56] 2024 3,554 full Audio, Lyrics Quadrants - S P
Popular Hooks [57] 2024 38,694 hooks Video, Audio, Lyrics Quadrants - S P
II Datasets
@@
Table I presents an extensive overview of emotion-annotated music datasets. These datasets vary in size, annotation granularity, and focus on either perceived or induced emotions. We have attempted to provide an exhaustive overview of emotion datasets. This was achieved by using Google Scholar with search terms such as ‘music emotion dataset’, ‘affective music dataset’, as well as following references within these articles. Before delving deeper into the datasets, it is essential to understand the emotion representations commonly referenced in the field.
@@
Categorical models include the Geneva Emotional Music Scales (GEMS) [27]. This model was specifically designed for music-induced emotions and consists of 45 emotion tags grouped into nine categories, including amazement, solemnity, tenderness, nostalgia, calmness, power, joyful activation, tension, and sadness. The datasets listed in Table I use a variety of emotion representation models, with Russell’s model being the most popular dimensional representation. Some datasets do not use a specific emotion model, for instance, the MTG-Jamendo dataset [36] consists of freely assigned tags by the listeners, resulting in a diverse and comprehensive set of 56 tags, spanning from ‘melancholic’ to ‘upbeat’. A subset of this dataset is used for the ‘Emotion and Theme Recognition in Music’ Task of MediaEval. In Table I, the number of tags used to represent the emotions are listed in the column ‘Categorical’.
@@
Induced versus perceived Emotion labels in the datasets can represent either perceived or induced emotions. Perceived emotions are emotions that listeners consciously recognize within the music itself. Induced emotions, on the other hand, are emotions that listeners experience as a result of listening to the music, which involve an actual emotional experience provoked by the stimulus. This emotional experience can be influenced by context, memories, and personal experiences, and may differ from the perceived emotion [62].
@@
Most of the datasets in Table I use perceived emotion labels, which are easier to annotate. For induced emotion labels, researchers have used psycho-physiological measurements ranging from electromyogram (EMG), volume pulse (BVP), electrocardiograms (ECG), skin conductance, respiration rate, heart rate, to electroencephalograms (EEG). There have been many studies in psychology that use such biosensors to explore how music can influence our emotions [63, 64, 65, 66, 67]. Since these studies include medical data, the datasets are not often public. However, there are a few datasets with music and its induced emotions. Firstly the DEAP dataset [22] includes EEG, facial video recordings, as well as peripheral physiological signals that were recorded while they watched music videos. In addition, they collected perceived emotion ratings in terms of arousal, valence, like/dislike, dominance and familiarity. Second, the HKU956 dataset [44] records five kinds of physiological signals (i.e., heart rate, electrodermal activity, blood volume pulse, inter-beat interval, and skin temperature) of participants as they listen to music, along with reported emotions in the arousal and valence dimensions. Third, the Music-Mouv’ dataset [52] investigates the impact of emotional context induced by music on gait initiation, focusing on anticipatory postural adjustments crucial for the elderly and individuals with Parkinson’s disease. It includes subjective emotional responses, physiological data from wristbands, and biomechanical data from shoe insoles equipped with sensors collected during and after music listening. Finally, SiTunes [55] includes physiological signals (i.e., heart rate, activity intensity, activity step, and activity type) measured both before and after music listening, alongside environmental data (i.e., time of day, weather information, and location) recorded during users’ daily lives. The MER models that can predict induced emotions have various medical applications, such as curating playlists to guide patients to different emotional states [5]. An interesting study by [68] concludes that music tends to induce the emotion that is perceived, enabling researchers to use perceived emotion datasets for developing emotion-inducing models.
@@
Modalities Music comes in many formats, the most common one being ‘audio’. Audio files can either be raw waveforms or compressed .mp3 files. However, we should not neglect the MIDI format, a popular format still often used by music producers, composers and performers. Whereas most datasets contain audio files, only a handful focus on MIDI: 1) the VGMIDI dataset [37], which contains continuous valence/arousal ratings for MIDI files of piano arrangements for video game soundtracks; 2) the Panda et al.’s dataset [23], which provides a diverse collection of audio clips, lyrics, and aligned MIDI files, contains 28 emotion labels grouped into five emotion clusters derived from a cluster analysis of online tags by [15]; and 3) the EMOPIA dataset [39], which contains paired piano music audio with MIDI that has emotion annotations of the four high/low valence/arousal quadrants.
@@
Many sources can elicit emotion. For instance, video, or lyrics may also affect our perceived or induced emotions. Some datasets offer alternative multimedia streams such as emotion-rated music videos from a variety of genres, including pop, rock, classical, and jazz as in the DEAP dataset [22]; or text of lyrics with the annotated music in the DMDD dataset [35]. The MuVi dataset [46] even offers isolated modality ratings for music videos. In this study, the raters were presented with either the music video, the music alone, or the muted video. The final dataset contains ratings for each of these modalities separately as well as together. This allows [46] to build a model on pure isolated modalities, which proved to be more accurate than a traditional model.
The RAVDESS dataset [34] is a multimodal dataset containing both speech and music with emotional expressions: calm, happy, sad, angry, fearful, surprise, and disgust for speech; and calm, happy, sad, angry, and fearful for music. It provides a rich combination of modalities, including Audio-only, Audio-Video, and Video-only formats.
Additionally, the EmoMV dataset [49] focuses on affective music-video correspondence learning by providing labeled pairs of music and video segments that either match or mismatch in terms of emotional content. This dataset enables models to learn the affective alignment between audio and visual modalities, making it particularly valuable for emotion-based matching and retrieval tasks.
Recently, the Popular Hooks dataset [57] was introduced as a multimodal dataset that contains 38,694 popular musical hooks (i.e., memorable sections of songs) with synchronized MIDI, music video, audio, and lyrics. It also offers detailed (predicted) labels for high-level musical attributes such as tonality, structure, genre, emotion, and region. Leveraging a pre-trained multimodal music emotion recognition framework, the dataset provides predicted emotion labels, which were evaluated through a user study.
Lastly, datasets like MERP [45] and EMOPIA [39] provide metadata on song features and listener demographics, which can be useful for detailed analysis and personalized emotion prediction.
@@
Dataset size Compared to affective datasets available in other domains (e.g., the Sentiment140 dataset [69] in NLP, which contains 1.6 million tweets annotated as positive or negative), the size of the available datasets is still very limited, with the largest dataset containing 109,269 instances. In addition to the number of instances in datasets, we also notice a difference in the length of the instances. A number of datasets (e.g. MERP [45], VGMIDI [37], and HKU956 [44]) offer ratings for full-length songs. In datasets with dynamic ratings throughout the song, this may provide a means for researchers to analyse how our emotions evolve throughout a song. Other datasets focus on short fragments, often ranging from 30 seconds to 1-minute fragments (e.g. DEAM [32], EMOPIA [39], and EMMA [54]), but even as short as 10s as is the case for the MER500 dataset which consists of Indian Hindi film music [40].
Variety of music The genres covered in these datasets vary widely, reflecting the diverse nature of music. For instance, the DEAP dataset [22] includes classical as well as jazz pieces, which are known for their rich emotional and structural complexities. The VGMIDI dataset on the other hand, [37] focuses exclusively on video game soundtracks, a genre that often aims to evoke specific emotions to enhance the gaming experience. Rock and electronic music can primarily be found in the DEAM dataset [32], both genres that are prevalent in contemporary music culture. This genre diversity allows researchers to select datasets that align with their specific research goals and to explore how different genres affect emotional perception and annotation. Additionally, studying a variety of genres can help in developing more robust and generalized MER models that can perform well across different types of music.
Annotation Process Annotation processes and label distribution vary significantly across datasets, impacting the effectiveness and reliability of machine learning models in MER. Some datasets, like MTG-Jamendo [36], rely on freely assigned tags by listeners, allowing for a broad range of emotional descriptors. In contrast, DEAP [22] and HKU956 [44] use a predetermined list of emotions to ensure consistency in emotion annotations. The number of annotators also differs, with some datasets involving large crowdsourced annotations, providing diverse perspectives, while others use smaller, more controlled groups to maintain uniformity. Additionally, the logic behind selecting music segments varies, with some datasets focusing on full-length songs and others on shorter clips to capture transient emotional responses.
Label distribution and splits The distribution of labels across datasets can be imbalanced, affecting the performance of machine learning models. For instance, the MTG-Jamendo dataset [36] contains a diverse set of 56 tags, but the frequency of these tags can vary significantly. In contrast, the DEAP dataset [22] provides a more balanced set of labels across arousal and valence dimensions. Finally, some datasets, such as DEAM [32], propose fixed splits for training and testing. This facilitates the replication of experiments and ensures consistent evaluation metrics which are easily used as a benchmark.
@@
In sum, there is a variety of emotion-annotated datasets available as shown in Table I. They differ in terms of emotion representation models, annotations labels, as well as music format and perceived versus induced emotion labels. Understanding these variations is crucial for developing robust and accurate MER models that can generalize across different datasets and musical genres. The challenges and future research directions related to datasets are discussed in detail in Section V.
III Evaluation Protocols
@@
Evaluation metrics play a crucial role in assessing the performance of MER systems. Depending on which type of emotion ratings are being predicted (dimensional versus categorical), the evaluation metrics change as we are dealing with a regression or a classification task respectively. Commonly used evaluation metrics for categorical MER systems include accuracy, precision, area under the ROC curve (AUC), and confusion matrices. In the case of regression MER models, metrics such as mean squared error (MSE), $R^{2}$, and Pearson correlation coefficient [70] are used. These metrics provide insights into the effectiveness of MER models when it comes to emotions represented by dimensional models.
@@
Evaluating MER systems goes beyond just defining a common metric. Due to the inherent differences between datasets, a comparison across datasets is often not possible. For an in-depth discussion on this, the reader is referred to Section V. Within one dataset, however, it is possible to establish benchmarks and compare the performance of different models, provided that the train/test split is shared. There are some initiatives to facilitate the comparison between models, such as competitions, as well as individual papers that offer clear data splits and metrics [71, 72, 73, 74, 75].
@@
Competitions and challenges provide valuable platforms for evaluating and comparing the performance of MER systems. These initiatives often involve standardized datasets, evaluation protocols, and metrics, enabling researchers to benchmark their algorithms against state-of-the-art methods. Existing benchmarking initiatives and competitions in MER include the ‘Audio K-POP Mood Classification’ task in MIREX (Music Information Retrieval Evaluation eXchange) (last organized in 2019) [^2] https://www.music-ir.org/mirex/wiki/2019:Audio_K-POP_Mood_Classification/ , the ‘Emotion in Music’ task in MediaEval (last organized in 2015) [^3] http://www.multimediaeval.org/mediaeval2015/emotioninmusic2015/, and the ‘Emotion and Theme Recognition in Music Using Jamendo’ in MediaEval (last organized in 2021) [^4] https://multimediaeval.github.io/2021-Emotion-and-Theme-Recognition\\-in-Music-Task/.
@@
TABLE II: Overview of Selected Music Emotion Recognition Models since 2020.
Ref. Year Modalities Approach Emotion Model Dataset
[76] 2020 Audio, Lyrics Machine learning (e.g., SVM, NB) Russell PMEmo
[77] 2020 Audio CNN, LSTM Russell Solymani et al.’s dataset
[78] 2020 Audio LSTM 4 classes Self-built
[79] 2020 Audio CNN-LSTM 4 classes Last.fm
[80] 2020 Audio Attentive LSTM Russell Solymani et al.’s dataset
[81] 2020 Audio Source separation, CNN Russell PMEmo
[82] 2020 Audio Cochleogram, CNN Russell Solymani et al.’s dataset
[83] 2020 Audio CNN, Local Attention Quadrants Soundtrack, Bi-Modal
[72] 2020 Audio Attention Based Neural Networks 56 classes MTG-Jamendo
[71] 2020 Audio Triplet Neural Networks Russell DEAM
[84] 2021 Audio CNN-LSTM Russell DEAM
[85] 2021 Audio CNN, LSTM+DNN 3 classes Self-built
[73] 2021 Audio Clustering-based Ensembles 56 classes MTG-Jamendo
[74] 2021 Audio Semi-Supervised, Noisy Student Training 56 classes MTG-Jamendo
[75] 2021 Audio Frequency Dependent Convolutions 56 classes MTG-Jamendo
[86] 2021 Audio CNN, Co-teaching Training Strategy 56 classes MTG-Jamendo
[87] 2021 Audio LSTM, Pretrained Models Russell Self-built
[88] 2021 Lyrics Transformers Russell MoodyLyrics, MER
[89] 2021 Audio Source-separation based explainer Russell DEAM, Midlevel, PMEmo
[90] 2021 Audio Generative Adversarial Network 2 classes Self-built
[91] 2021 Audio, Video CNN 6 classes Self-built
[92] 2021 Audio Linear Regressors Russell Self-built
[93] 2022 Audio Clustering, Machine Learning (e.g., SVM) Russell Solymani et al.’s dataset
[94] 2022 MIDI Multi-Task Learning Quadrants EMOPIA, VGMIDI
[95] 2022 Audio Feature Selection, SVR, RF Russell DEAM
[46] 2022 Audio, Video LSTM Russell Self-built
[96] 2022 Lyrics Deep Learning, BERT Quadrants MIR Lyrics Emotion
[97] 2022 Audio CNN, LSTM Russell PMEmo
[98] 2022 Audio SVM, Random Forest, MLP Russell DEAM
[99] 2023 Audio, Lyrics Multi-Modality Russell DMDD
[100] 2023 Audio Inception-GRU Residual Quadrants Soundtrack
[101] 2023 Lyrics State Space Models 6 classes LyricsEmotions
[102] 2023 Lyrics Fine-tuned XLMRoBERTa Russell Self-built
[50] 2023 Audio, Lyrics CNN-LSTM, XLNet Transformers 3 classes Self-built
[103] 2023 Audio Attention Mechanism Russell DEAM, PMEmo
[104] 2023 Audio CNN 4 classes Musical Excerpts
[105] 2023 Audio, Lyrics Transformers Quadrants MoodyLyrics
[106] 2024 Audio RNN, BRNN, LSTM Quadrants 4Q audio, MTG-Jamendo
[107] 2024 Audio, Lyrics VGGish, ALBERT 4 classes DEAM, FMA
[108] 2024 Audio, Lyrics Stacked Ensemble Models 4 classes Self-built
[109] 2024 Audio Multi-scale Parallel Convolution 8 classes PMEmo, Soundtrack, RAVDESS
[110] 2024 Audio CNN with Differential Evolution 4 classes self-built, DEAM
IV Models and approaches
In this section, we briefly touch upon some of the more recent MER models, highlighting the state-of-the-art approaches over the last few years. This is not an exhaustive overview, and for a more comprehensive review, readers are referred to other survey papers [11, 12, 13, 9, 14]. The aim of this section is to point out the current state-of-the-art and various approaches over the last five years. The models were selected through Google Scholar searches using search terms ‘music emotion prediction model’, ‘affective music prediction model’, and ‘music emotion recognition model’ from dates 2020, as well as by following links in the articles.
Table II presents an overview of selected MER models released since 2020, summarizing their modalities, approaches, emotion models, and datasets. This table highlights the diversity in methodologies and datasets used in recent research.